### v1

@@ Abstract
Molecular Dynamics (MD) simulations are vital for exploring complex systems in computational physics and chemistry. While machine learning methods dramatically reduce computational costs relative to ab initio methods, their accuracy in long-lasting simulations remains limited. Here we propose dynamic training (DT), a method designed to enhance model performance over extended MD simulations. Applying DT to an equivariant graph neural network (EGNN) on the challenging system of a hydrogen molecule interacting with a palladium cluster anchored to a graphene vacancy demonstrates a superior prediction accuracy compared to conventional approaches. Crucially, the DT architecture-independent design ensures its applicability across diverse machine learning potentials, making it a practical tool for advancing MD simulations.
@@ 1 Introduction
Molecular dynamics (MD) simulations have proven to be a powerful and versatile tool, providing valuable insights into the mechanisms behind complex phenomena [1, 2, 3]. Moreover, MD simulations also hold significant promise in optimizing and accelerating discovery of novel materials [4]. However, with current MD approaches, one often needs to choose between accuracy and simulation speed. Ab initio methods such as density functional theory (DFT) provide highly accurate predictions but are computationally expensive and scale poorly with system size. In contrast, classical force fields offer near-linear scaling with system size, but often lack accuracy and the transferability required for application to diverse systems. In recent years machine learning potentials (MLPs) have emerged as a powerful alternative to the aforementioned methods as they offer the possibility of achieving accuracy comparable to that of ab initio methods, while maintaining near-linear scaling with system size due to their predominantly local nature. MLPs come in various forms, including kernel methods [5, 6, 7], permutationally invariant polynomials [8, 9], and neural networks [10, 11, 12, 13, 14, 15, 16, 17, 18]. Among these, neural network potentials (NNPs) have emerged as a particularly promising avenue for creation of accurate and efficient multidimensional potential energy surfaces (PES) [19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. However, the application of neural network potentials is not without challenges. The requirement for extensive training data can be a substantial barrier, particularly for systems for which high-quality reference calculations are computationally expensive. Furthermore, the standard training paradigm, which focuses on minimizing the global error in single-step predictions, may not adequately capture local intricacies present in the PES. This disparity becomes apparent in applications such as MD simulations, where the cumulative effect of errors and exposure to varying temperatures can drive the system into regions where the potential is less accurately learned, frequently causing instabilities in the simulated dynamics [29].
@@
In this work, we propose a dynamic training (DT) approach for enhancing the training of NNPs from ab-initio molecular dynamics (AIMD) simulations. We apply this strategy to an equivariant graph neural network (EGNN), resulting in what we term DT-EGNN. In contrast to conventional approaches that process data points in isolation, our method explicitly accounts for the sequential nature of MD simulations by including integration of equations of motion into the training process of a neural network. Thus, it enables direct comparison between predicted simulations and AIMD reference data, enhancing the ability of a model to capture the temporal evolution of the system. To exemplify this statement, a comprehensive and challenging dataset of AIMD simulations describing the dynamics of molecules interacting with clusters anchored in graphene vacancies (H2/Pd6@G) [30] is used and we demonstrate that DT-EGNN achieves higher accuracy compared to conventional training methods while indicating promising data efficiency.
@@ 2.1 Method Description
Development of machine learning models in computational chemistry is often hindered by the scarcity and high computational cost of accurate data. One way to speed up the process of creating a dataset is to use AIMD. Many widely-used datasets, including MD17 [6], Open Catalyst Project [31], and ANI-1x [32] are either partially or fully generated from AIMD. The common way of utilizing these datasets is to randomly select atomic structures for training, validation, and testing. While this approach simplifies data handling, it discards valuable temporal information present in the simulations. Given that many NNP applications revolve around performing MD simulations, it is reasonable to expect that incorporating temporal structure within the training process would enhance the NNP quality. Therefore, in this work, we propose treating each data point as a subsequence of an AIMD simulation rather than as an isolated atomic configuration. This in turn enables us to incorporate molecular dynamics directly into the training process of a NNP. In order to implement this approach, we have chosen an EGNN as our NNP architecture (see Fig. 1a). Detailed information about the EGNN architecture employed in this work is provided in the Methods section.
@@
In order to perform stable dynamics within the forward pass, we must ensure that the predictions on initial atomic structures are as accurate as possible. To achieve this, the training process starts with the standard practice of minimizing prediction errors of energies and forces on single atomic structures, which correspond to the initial atomic structure of each training point. It is useful to frame them as subsequences of length one (). Once the convergence criteria is met, instead of terminating the training process, we continue with the training by incrementing the subsequence length by one. Consequently, this makes the training iteration to consist of more parts as summarized in Fig. 1b. Similarly to , it starts with the prediction of the energy and forces for the initial atomic structures. These forces, in conjunction with velocities and positions, which were stored in the data preprocessing part, are utilized to derive the next-step atomic coordinates via the Euler method. These coordinates can now be used to build a graph in a similar way as described before. Upon generation of a new graph, the model predicts new atomic forces, which are then combined with the forces from the previous step to update the velocities using the velocity Verlet algorithm. This forms a loop that occurs once if the subsequence length is two, and more generally times if the length of the subsequence is . Once the model yields the predictions of energies and atomic forces for the whole subsequence, they are compared to those obtained in the corresponding AIMD calculations that were recorded during data preprocessing step. The loss function, as well as other important details regarding the training process are described in the Methods section.
@@
Here, a remark regarding the calculation of the atomic neighborhoods in DT is in order. For subsequence lengths greater than one, each update of the atomic positions during the dynamical training can modify the underlying atomic neighborhood structures and, therefore, they must be recalculated. The typical method used to construct a neighborhood in an atomistic system is the radius graph, where neighbors comprise all atoms within a sphere of radius R from a given atom. However, this approach presents a challenge due to the nature of the greater-than-or-equal-to function, which acts as a step function in determining neighbor status. Since the step function is not differentiable, the computational graph would be disconnected leading to poor learning. While employing some kind of smoother function (e.g., sigmoid function) might seem like a natural solution, it remains unclear how to implement message passing on a continuous scale in this context without taking all atoms as neighbors. This in turn would be problematic both in terms of scaling the system and applying the methodology to periodic systems because the computational cost would be prohibitively expensive. To address this issue, we treat the atomic neighborhood as a parameter derived from the corresponding AIMD calculations (see Supplementary Note 1). Importantly, this problem only exists during the training phase as we do not require differentiability during inference.
@@ 2.2 DT accuracy and efficiency
We validate the proposed DT method on a challenging dataset of AIMD simulations. The dataset contains 100 DFT-based microcanonical trajectories integrated with a timestep fs, resulting in a total of 228,925 atomic configurations. In these simulations, a H2 molecule with an initial translational energy of 0.125 eV interacts with a substrate equilibrated at 300 K. The substrate consists of a Pd6 cluster anchored to a vacancy of a graphene layer. Different processes such as H2 scattering after one or multiple bouncing events, as well as H2 adsorption and H2 dissociation on the Pd6 cluster that often involve Pd6 isomerization are observed [30]. Using these data, we have trained several DT-EGNN models that basically differ in the size of the employed training sets (number of training points and values). All these models served us to test and confirm the accuracy and efficiency of the DT method as follows. First we show that increasing the subsequence length during the training process increases the accuracy of the DT-EGNN models. Next, we demonstrate that increasing the subsequence length also reduces the errors in atomic configurations not seen during the training process, underscoring the data efficiency of the method.
@@
Common practices often involve using validation points that closely resemble training points. In this work however, we adopt a different strategy. Our training set consists of 226,825 points and each training point has a maximum subsequence length , whereas the validation set contains 2048 points, each having a subsequence length . In other words, each time we evaluate our model against the validation set we perform 2048 MD simulations for 60 integration steps. It is important to point out that while the subsequence length of a training point changes during the training process every time a convergence is reached (see Fig. 1b), it stays fixed for a validation point. This provides us with a constant benchmark throughout the training process, while also giving us the opportunity to guide the training process towards a model that best aligns with the demands of a long-lasting molecular dynamics simulation. The validation error curve during the training process of one of the calculated DT-EGNN models for H2/Pd6@G is shown in Fig. 2. The employed error metric, defined in equation (13) in the Methods section and abbreviated as MEPA, represents the mean error of atomic forces per atom type and per simulation step. The validation curve exhibits the typical steep decrease that is obtained at the beginning of the training process. The scheduler decreases the learning rate when the validation error shows no improvement over a patience period. When the patience period is exceeded at the minimum learning rate, we consider the model converged for the current subsequence length. First such convergence is marked by the red dashed line in Fig. 2. Once this criterium is met, the subsequence length is increased by one, and the learning rate reset. This reset value turned out to be a very important hyperparameter because a too small value might cause the neural network to stay trapped in the local minimum it found itself after the last convergence, whereas a too large value might destabilize the learning process. Details of all hyperparameters used in the training processes are provided in the Methods section. The novel ingredient in the DT method is the information on the system dynamics that is incorporated through the subsequence length hyperparameter. The benefit of adding such information is confirmed in Fig. 2, in which we observe that the validation error decreases with the increase in the subsequence length. However, this decrease is not strictly monotonic due to the fluctuations which largely align with the increase in learning rate following convergence. Since the validation error represents the average error per simulation step, these validation error reductions have an amplified significance. This is because the accumulation of errors at each step becomes increasingly important over longer time scales.
@@
A key aspect of any machine learning method is its efficiency in utilizing training data. This is especially true for atomistic systems where ab-initio data are scarce and expensive to come by. We have investigated how well our approach generalizes to atomic structures unseen in the training set. Leaving the model architecture unchanged, we have selected for training a subset of 50 simulations from the original 100 AIMD simulations. This amounts to 108,019 training points with the maximum subsequence length set to = 15. The validation set consists of 512 points with a subsequence length equal to . The 512 points used for validation are chosen from the same subset of 50 simulations used for training so that the other half of the AIMD simulations remain completely invisible to the model. In order to examine how the training subsequence length affects the model performance, we save the model with the lowest validation error at each subsequence length during the training process. Thus, as = 15, we end up with fifteen models with each model representing the best performing model for its respective training subsequence length. Using a set of 50 previously unseen AIMD simulations, we randomly sampled 256 atomic structures. Each structure served as a starting point for a 120-timesteps simulation, which is then compared to the reference AIMD data. The resulting mean absolute errors (MAE) in atomic forces for each of the fifteen models are shown in Fig. 3. We observe a consistent decrease in the prediction error with increasing subsequence length, despite the number of unique atomic structures in the training set staying constant throughout the training process.
@@ 2.3 Comparison to conventional training methods
So far we have examined the properties and performance of the DT approach in isolation. However, to establish its practical value, we compare it to conventional NNP training methods. We evaluate the DT-EGNN model trained using the complete AIMD data set with and , against the state-of-the-art MACE NNP and two additional NNPs (EGNN-MSE and EGNN-MEPA) trained using traditional single-point configurations for both training and validation. The latter two NNPs share identical architecture and number of parameters with our DT-EGNN model. However, EGNN-MSE has the mean squared error (MSE) of energies and forces as a loss function (see equation (16)), whereas EGNN-MEPA has the same loss function as DT-EGNN at subsequence length (equations (9) to (11)). All NNPs were trained on the complete dataset of 100 AIMD simulations as detailed in the Methods section. To test the performance beyond the training simulation length of the DT-EGNN models, we selected one random atomic structure from each of the 100 trajectories, resulting in 100 test configurations. From each test configuration and for each NNP, we performed 300-timesteps MD simulations using the AIMD timestep of 0.5 fs. The atomic forces obtained at each integration step are compared to their AIMD counterparts. The corresponding MEPA values obtained with the different NNPs are summarized in Table 1 and also plotted in Fig. 4 to highlight the DT-EGNN performance with the training subsequence length. With all EGNN models being structurally equal, it is interesting to compare how different training design choices impact the performance. Since EGNN-MEPA has the same loss function as DT-EGNN at training subsequence length equal to one, we would expect it to perform the same as the DT-EGNN model at . However, there is a difference in the MEPA value of 2 meV/Å that comes from the DT-EGNN validation scheme using simulations of length 60, which makes it more optimized for this task. Notably, EGNN-MSE performs the poorest. This is a consequence of having an unbalanced number of atoms between the different atoms types in this system, namely, 49 carbon atoms, six palladium atoms and only two hydrogen atoms. Minimizing the mean squared error or the mean absolute error metrics pushes the network to minimize the error on the carbon atoms, as that is the best strategy to minimize the overall mean error of a given configuration (see Supplementary Note 2). This problem is further exacerbated by the fact that hydrogen atoms cover the biggest phase space volume in our system. However, if we first take the mean of the error per each atom type and take the sum of these errors, we have a much more realistic measure of how the network is performing. Surprisingly, DT-EGNN is outperforming the state-of-the-art MACE model already at subsequence length equal to one and, as shown in Fig. 4, it continues to improve its performance as the training subsequence length increases.
@@
In conclusion, we have introduced dynamic training, a unique methodology to train NNPs optimized for long-lasting molecular dynamics. Our approach extends beyond the conventional single-point training paradigm by incorporating sequences of consecutive atomic configurations that provide information on the system dynamics. We have validated the method on a challenging dataset of 100 ab-initio molecular dynamics simulations of impinging on a substrate consisting of a Pd6 cluster anchored to a graphene vacancy. With the help of this dataset we have explored the relation between the accuracy of a model and the training subsequence length. We found that increasing the subsequence length during the training process does indeed improve the accuracy of the DT-EGNN. Furthermore, we have demonstrated the ability of the DT-EGNN method to generalize on atomic structures not present in the training set. Finally, comparing the DT-EGNN to conventionally trained NNPs on simulation lengths much larger than those present in the training set, we show that DT-EGNN outperforms the EGNNs with identical architecture but trained in conventional fashion. Moreover, we also show that it performs better than the state-of-the-art model MACE.

### v2

@@ Abstract
Molecular Dynamics (MD) simulations are vital for exploring complex systems in computational physics and chemistry. While machine learning methods dramatically reduce computational costs relative to ab initio methods, their accuracy in long-lasting simulations remains limited. Here we propose dynamic training (DT), a method designed to enhance accuracy of a model over extended MD simulations. Applying DT to an equivariant graph neural network (EGNN) on the challenging system of a hydrogen molecule interacting with a palladium cluster anchored to a graphene vacancy demonstrates a superior prediction accuracy compared to conventional approaches. Crucially, the DT architecture-independent design ensures its applicability across diverse machine learning potentials, making it a practical tool for advancing MD simulations.
keywords:
[…]
@@ 1 Introduction
Molecular dynamics (MD) simulations have proven to be a powerful and versatile tool, providing valuable insights into the mechanisms behind complex phenomena 1, 2, 3. Moreover, MD simulations also hold significant promise in optimizing and accelerating discovery of novel materials 4. However, with current MD approaches, one often needs to choose between accuracy and simulation speed. Ab initio methods such as density functional theory (DFT) provide highly accurate predictions but are computationally expensive and scale poorly with system size. In contrast, classical force fields offer near-linear scaling with system size, but often lack accuracy and the transferability required for application to diverse systems. In recent years machine learning potentials (MLPs) have emerged as a powerful alternative to the aforementioned methods as they offer the possibility of achieving accuracy comparable to that of ab initio methods, while maintaining near-linear scaling with system size due to their predominantly local nature. MLPs come in various forms, including kernel methods 5, 6, 7, permutationally invariant polynomials 8, 9, and neural networks 10, 11, 12, 13, 14, 15, 16, 17, 18. Among these, neural network potentials (NNPs) have emerged as a particularly promising avenue for creation of accurate and efficient multidimensional potential energy surfaces (PES) 19, 20, 21, 22, 23, 24, 25, 26, 27, 28. However, the application of neural network potentials is not without challenges. The requirement for extensive training data can be a substantial barrier, particularly for systems for which high-quality reference calculations are computationally expensive. Furthermore, the standard training paradigm, which focuses on minimizing the global error in single-step predictions, may not adequately capture local intricacies present in the PES. This disparity becomes apparent in applications such as MD simulations, where the cumulative effect of errors and exposure to varying temperatures can drive the system into regions where the potential is less accurately learned, frequently causing instabilities in the simulated dynamics 29.
@@
In this work, we propose a dynamic training (DT) approach for enhancing the training of NNPs from ab-initio molecular dynamics (AIMD) simulations. We apply this strategy to an equivariant graph neural network (EGNN), resulting in what we term DT-EGNN. In contrast to conventional approaches that process data points in isolation, our method explicitly accounts for the sequential nature of MD simulations by including integration of equations of motion into the training process of a neural network. Thus, it enables direct comparison between predicted simulations and AIMD reference data, enhancing the ability of a model to capture the temporal evolution of the system. To exemplify this statement, a comprehensive and challenging dataset of AIMD simulations describing the dynamics of molecules interacting with clusters anchored in graphene vacancies (H2/Pd6@G) 30 is used and we demonstrate that DT-EGNN achieves higher accuracy compared to conventional training methods while indicating promising data efficiency.
@@ 2.1 Method Description
Development of machine learning models in computational chemistry is often hindered by the scarcity and high computational cost of accurate data. One way to speed up the process of creating a dataset is to use AIMD. Many widely-used datasets, including MD17 6, Open Catalyst Project 31, and ANI-1x 32 are either partially or fully generated from AIMD. The common way of utilizing these datasets is to randomly select atomic structures for training, validation, and testing. While this approach simplifies data handling, it discards valuable temporal information present in the simulations. Given that many NNP applications revolve around performing MD simulations, it is reasonable to expect that incorporating temporal structure within the training process would enhance the NNP quality. Therefore, in this work, we propose treating each data point as a subsequence of an AIMD simulation rather than as an isolated atomic configuration. This in turn enables us to incorporate molecular dynamics directly into the training process of a NNP. In order to implement this approach, we have chosen an EGNN as our NNP architecture (see Figure 1a). Detailed information about the EGNN architecture employed in this work is provided in the Methods section.
@@
In order to perform stable dynamics within the forward pass, we must ensure that the predictions on initial atomic structures are as accurate as possible. To achieve this, the training process starts with the standard practice of minimizing prediction errors of energies and forces on single atomic structures, which correspond to the initial atomic structure of each training point. It is useful to frame them as subsequences of length one (). Once the convergence criteria is met, instead of terminating the training process, we continue with the training by incrementing the subsequence length by one. Consequently, this makes the training iteration to consist of more parts as summarized in Figure 1b. Similarly to , it starts with the prediction of the energy and forces for the initial atomic structures. These forces, in conjunction with velocities and positions, which were stored in the data preprocessing part, are utilized to derive the next-step atomic coordinates. These coordinates can now be used to build a graph in a similar way as described before. Upon generation of a new graph, the model predicts new atomic forces, which are then combined with the forces from the previous step to update the velocities using the velocity Verlet algorithm. This forms a loop that occurs once if the subsequence length is two, and more generally times if the length of the subsequence is . Once the model yields the predictions of energies and atomic forces for the whole subsequence, they are compared to those obtained in the corresponding AIMD calculations that were recorded during data preprocessing step. The loss function, as well as other important details regarding the training process are described in the Methods section.
@@
Here, a remark regarding the calculation of the atomic neighborhoods in DT is in order. For subsequence lengths greater than one, each update of the atomic positions during the dynamical training can modify the underlying atomic neighborhood structures and, therefore, must be recalculated. The typical method used to construct a neighborhood in an atomistic system is the radius graph, where neighbors comprise all atoms within a sphere of radius R from a given atom. However, this approach presents a challenge due to the nature of the greater-than-or-equal-to function, which acts as a step function in determining neighbor status. Since the step function is not differentiable, the computational graph would be disconnected leading to poor learning. While employing some kind of smoother function (e.g., sigmoid function) might seem like a natural solution, it remains unclear how to implement message passing on a continuous scale in this context without taking all atoms as neighbors. This in turn would be problematic both in terms of scaling the system and applying the methodology to periodic systems because the computational cost would be prohibitively expensive. To address this issue, we treat the atomic neighborhood as a parameter derived from the corresponding AIMD calculations (see Supplementary Note 1). Importantly, this problem only exists during the training phase as we do not require differentiability during inference.
@@ 2.2 DT accuracy and efficiency
We validate the proposed DT method on a challenging dataset of AIMD simulations. The dataset contains 100 DFT-based microcanonical trajectories integrated with a timestep fs, resulting in a total of 228,925 atomic configurations. In these simulations, a H2 molecule with an initial translational energy of 0.125 eV interacts with a substrate equilibrated at 300 K. The substrate consists of a Pd6 cluster anchored to a vacancy of a graphene layer. Different processes such as H2 scattering after one or multiple bouncing events, as well as H2 adsorption and H2 dissociation on the Pd6 cluster that often involve Pd6 isomerization are observed 30. Using these data, we have trained several DT-EGNN models that basically differ in the size of the employed training sets (number of training points and values). All these models served us to test and confirm the accuracy and efficiency of the DT method as follows. First we show that increasing the subsequence length during the training process increases the accuracy of the DT-EGNN models. Next, we demonstrate that increasing the subsequence length also reduces the errors in atomic configurations not seen during the training process, underscoring the data efficiency of the method.
@@
Common practices often involve using validation points that closely resemble training points. In this work however, we adopt a different strategy. Our training set consists of 226,825 points and each training point has a maximum subsequence length , whereas the validation set contains 2048 points, each having a subsequence length . In other words, each time we evaluate our model against the validation set we perform 2048 MD simulations for 60 integration steps. It is important to point out that while the subsequence length of a training point changes during the training process every time a convergence is reached (see Figure 1b), it stays fixed for a validation point. This provides us with a constant benchmark throughout the training process, while also giving us the opportunity to guide the training process towards a model that best aligns with the demands of a long-lasting molecular dynamics simulation. The validation error curve during the training process of one of the calculated DT-EGNN models for H2/Pd6@G is shown in Figure 2. The employed error metric, defined in eq (13) in the Methods section and abbreviated as MEPA, represents the mean error of atomic forces per atom type and per simulation step. The validation curve exhibits the typical steep decrease that is obtained at the beginning of the training process. The scheduler decreases the learning rate when the validation error shows no improvement over a patience period. When the patience period is exceeded at the minimum learning rate, we consider the model converged for the current subsequence length. First such convergence is marked by the red dashed line in Figure 2. Once this criterium is met, the subsequence length is increased by one, and the learning rate reset. This reset value turned out to be a very important hyperparameter because a too small value might cause the neural network to stay trapped in the local minimum it found itself after the last convergence, whereas a too large value might destabilize the learning process. Details of all hyperparameters used in the training processes are provided in the Methods section. The novel ingredient in the DT method is the information on the system dynamics that is incorporated through the subsequence length hyperparameter. The benefit of adding such information is confirmed in Figure 2, in which we observe that the validation error decreases with the increase in the subsequence length. However, this decrease is not strictly monotonic due to the fluctuations which largely align with the increase in learning rate following convergence. Since the validation error represents the average error per simulation step, these validation error reductions have an amplified significance.
@@
A key aspect of any machine learning method is its efficiency in utilizing training data. This is especially true for atomistic systems where ab-initio data are scarce and expensive to come by. We have investigated how well our approach generalizes to atomic structures unseen in the training set. Leaving the model architecture unchanged, we have selected for training a subset of 50 simulations from the original 100 AIMD simulations. This amounts to 108,019 training points with the maximum subsequence length set to = 15. The validation set consists of 512 points with a subsequence length equal to . The 512 points used for validation are chosen from the same subset of 50 simulations used for training so that the other half of the AIMD simulations remain completely invisible to the model. In order to examine how the training subsequence length affects the model performance, we save the model with the lowest validation error at each subsequence length during the training process. Thus, as = 15, we end up with fifteen models with each model representing the best performing model for its respective training subsequence length. Using a set of 50 previously unseen AIMD simulations, we randomly sampled 256 atomic structures. Each structure served as a starting point for a 120-timesteps simulation, which is then compared to the reference AIMD data. The resulting mean absolute errors (MAE) in atomic forces for each of the fifteen models are shown in Figure 3. We observe a consistent decrease in the prediction error with increasing subsequence length, despite the number of unique atomic structures in the training set staying constant throughout the training process.
@@ 2.3 Comparison to conventional training methods
So far we have examined the properties and performance of the DT approach in isolation. However, to establish its practical value, we compare it to conventional NNP training methods. We evaluate the DT-EGNN model trained using the complete AIMD data set with and against two NNPs, namely MACE 18 and EGNN-MEPA, trained on a complete AIMD data set but in a conventional fashion. These two NNPs provide valuable benchmarks for comparison. EGNN-MEPA has identical architecture and number of parameters to DT-EGNN, differing only in the training method, while MACE represents a powerful, widely-used equivariant graph neural network with a computationally heavier architecture. To test performance beyond the training simulation length of the DT-EGNN models, we selected one random atomic structure from each of the 100 trajectories, resulting in 100 test configurations. From each test configuration and for each NNP, we performed 300-timesteps MD simulations using the AIMD timestep of 0.5 fs. The atomic forces obtained at each integration step are compared to their AIMD counterparts and the corresponding MEPA values are shown in Figure 4. As expected, DT-EGNN at subsequence length equal to one and EGNN-MEPA perform similarly since the only difference at that subsequence length between these two NNPs is the validation scheme of DT-EGNN. However, increasing the subsequence length leads to a clear reduction of MEPA values, demonstrating the benefits of DT even for molecular simulations significantly longer than those encountered during training. Moreover, DT-EGNN outperforms the state-of-the-art NNP MACE, indicating an important trade-off in NNP design. Complex architectures like MACE promise better extrapolation capabilities through built-in symmetry constraints. However, this advantage comes with a high computational cost during inference. Since inference time typically constitutes the larger share of computational expenses compared to training time, DT offers a compelling alternative. It acts as a regularization technique imposed by temporal correlations present in molecular dynamics data and increases an accuracy of a NNP without increasing the complexity of a NNP itself. It might, however, seem surprising that EGNN-MEPA and DT-EGNN (at subsequence length equal to one) have better performance than MACE. This can, at least partly, be explained in terms of the choice of loss function. By default, MACE uses the weighted mean squared error (MSE) of energies and forces as a loss function (eq (14)). However, such a loss function underperforms when systems contain disproportionate numbers of atoms across different atom types (see Supplementary Note 2). This problem is further exacerbated because hydrogen atoms, despite being the least abundant species with only two atoms present, cover the largest phase space volume in our system. Results of all models are summarized in Table 1.
@@
3 Conclusion
In conclusion, we have introduced dynamic training, a unique methodology to increase accuracy of NNPs for long-lasting molecular dynamics. Our approach extends beyond the conventional single-point training paradigm by incorporating sequences of consecutive atomic configurations that provide information on the system dynamics. We have validated the method on a challenging dataset of 100 ab-initio molecular dynamics simulations of impinging on a substrate consisting of a Pd6 cluster anchored to a graphene vacancy. With the help of this dataset we have explored the relation between the accuracy of a model and the training subsequence length. We found that increasing the subsequence length during the training process does indeed improve the accuracy of the DT-EGNN. Furthermore, we have demonstrated the ability of the DT-EGNN method to generalize on atomic structures not present in the training set. Finally, comparing the DT-EGNN to conventionally trained NNPs on simulation lengths much larger than those present in the training set, we show that DT-EGNN outperforms the EGNNs with identical architecture but trained in conventional fashion. Moreover, comparison with state-of-the-art NNP MACE reveals a significant advantage of DT. It enhances NNP accuracy without increasing architectural complexity, offering a computationally efficient alternative to complex architectures that impose higher inference costs. This demonstrates that leveraging temporal correlations through DT provides a practical pathway to improving molecular dynamics simulations while maintaining favorable computational performance.